OpenStudy (anonymous):

What is the greatest common factor of 12 and 20?

OpenStudy (anonymous):

4

OpenStudy (swissgirl):

The factors of 12: 1*12=12 2*6=12 3*4=12 Factors of 12=1,2,3,4,6,12

OpenStudy (swissgirl):

Factors of 20: 1*20=20 2*10=20 4*5=20 Factors of 20=1,2,4,5,10,20

OpenStudy (swissgirl):

So what is the greatest factor that both 20 and 12 have?

OpenStudy (swissgirl):

AND ITS 4